# Landcover classification and change detection using remote sensing and machine learning: a case study of Western Fiji

### v1

@@ Abstract
[inst1]Transitional Artificial Intelligence Research Group, School of Mathematics and Statistics,, UNSW Sydney, Sydney, Australia
@@
[…]
[inst3]Centre for AI and Innovation, Pingla Institute,, Sydney, Australia
@@ 2.6.2 Technical setup
In this study, we used 70% of the samples for training and 30% for tests. The model was trained yearly, and the training datasets were generated each year.
X
@@ 3.1 Selection of sample size
sample size train size test size accuracy precision recall f1 score
490 343 147 0.96599 0.96779 0.96599 0.96577
700 490 210 0.95238 0.95305 0.95238 0.95210
1050 735 315 0.98095 0.98123 0.98095 0.98095
1400 980 420 0.96190 0.96242 0.96190 0.96192
1750 1225 525 0.99048 0.99060 0.99048 0.99048
@@ 3.2 Model construction
Random forest: The hyper-parameters used in the random forest algorithm are selected using a five-stage grid search, including searching ‘max depth’, ‘max features’, ‘min samples leaf’, ‘min samples split’ and ‘n estimators’. The values of each parameter used in the experiment are stated in the table.
Artificial Neural Networks: The model’s training inputs are the 1225 vectors that were flattened from the image chips, where each vector has a size of 9*9*10.The parameters used in the Artificial Neural Networks were selected by the Grid Search method, including searching for the number of layers and nodes in the layer, finding an appropriate optimizer, and an activation function. Our goal of training is to improve the accuracy of classification. With this, the ’Accuracy’ is monitored. During the search, a check was done after each epoch. We allowed the process to be early-stopped when there was no improvement after ten epochs. We used 100 epochs and 32 batch size in training MLP, which was finally constructed with one hidden layer and 32 hidden neurons, using the ’Adam’ optimizer and rectifier linear units (Relu) activation function. We used the ’categorical cross-entropy’ function to measure the loss. The experiments were run ten times. With each run, the average accuracy was calculated based on 10-fold cross-validation. The experiment that gave the best average accuracy became our result.
Convolutional Neural Networks: The model’s training inputs are the the 1225 chips that were mapped with ground truth samples, where each chip had a size of 9*9 pixels and ten channels. We used three convolutional layers and located 32, 48, and 64 kernels in the layers. Each kernel has a size of 1*1 to help reduce the dimension. While training, we allowed 0.25 and 0.5 dropouts to avoid overfitting. The output from the convolutional layers was flattened and passed to the fully-connected layer. The activation function ’ReLu’ and the optimizer ’Adam’ were used. We used the ’sparse categorical cross-entropy function to measure the loss. The experiments were run ten times. With each run, the average accuracy was calculated based on 10-fold cross-validation. The experiment that gave the best average accuracy became our results.
@@ 3.3 Score comparison
In order to choose the model with the best classification performance, we have compared the accuracy scores of three supervised learning models, which were delivered by dividing the number of correctly classified samples by the total test samples. The classification accuracies for all three models were high. ANN had an accuracy of 95.24%, and CNN had an accuracy of 99.05%. The confusion matrices are shown in the Figure. In general, all three models performed well in classifying the Coastal Areas and wetlands while making mistakes in classifying the Grass/Agricultural Land, Forest and Water Body. We especially cared about the performance of the models in classifying the urban areas. The distribution of houses is sparse in the eastern side of our study area, close to the mountain areas, making distinguishing between bare land and urban area difficult due to their similar spectral reflectance. CNN performed best among the three models. This could be a good sign for us since an accurate classification of the urban areas is essential for change detection.
Table 2: Accuracy over 30 experimental runs for supervised learning models
Run CNN RF ANN
1 0.990476 0.986667 0.977143
2 0.984762 0.986667 0.984762
3 0.992381 0.990476 0.973333
4 0.992381 0.984762 0.975238
5 0.988571 0.990476 0.980952
6 0.990476 0.982857 0.980952
7 0.986667 0.988571 0.973333
8 0.984762 0.988571 0.984762
9 0.992381 0.986667 0.975238
10 0.994286 0.982857 0.975238
11 0.984762 0.988571 0.977143
12 0.992381 0.984762 0.980952
13 0.992381 0.986667 0.979048
14 0.994286 0.984762 0.967619
15 0.990476 0.998095 0.973333
16 0.990476 0.984762 0.975238
17 0.990476 0.988571 0.980952
18 0.986667 0.992381 0.982857
19 0.984762 0.982857 0.977143
20 0.992381 0.988571 0.980952
21 0.979048 0.994286 0.980952
22 0.994286 0.990476 0.980952
23 0.986667 0.982857 0.982857
24 0.988571 0.986667 0.977143
25 0.996190 0.982857 0.980952
26 0.980952 0.988571 0.982857
27 0.988571 0.988571 0.973333
28 0.990476 0.988571 0.980952
29 0.994286 0.994286 0.977143
30 0.990476 0.992381 0.982857
Mean 0.989524 0.987936 0.978539
Std 0.004117 0.003820 0.004092
@@
(a) ANN
(b) CNN
Figure 6: The ROC AUC curves of the deep learning models
@@ 3.5 Landcover Maps
(a) 2013
(b) 2014
(c) 2015
(d) 2016
(e) 2017
(f) 2018
(g) 2019
(h) 2020
(i) 2021
(j) 2022
(k) 2023
(l) 2024
Figure 8: xxxxx
@@ 3.6 Urban Area change detection
Based on the land cover maps that were generated by CNN, we created an urban expansion map and a replacement map for visualising the change in the urban area. The urban expansion map (Figure 3.6) shows how the urban area was expanded from 2013 to 2024. In the image, the non-urban area is represented by the white color. The replacement map (Figure 3.6) was generated by subtracting the 2013 land cover map from the 2024 land cover map. The background includes the area that was classified as urban in 2013 and other regions that remain non-urban.
The urban expansion map reveals that the urban area in our study has expanded outward along the existing periphery. The replacement map reveals that the growth of the urban area in the Nadi downtown, near the coastal line, corresponds to a substantial reduction in grass and agricultural Land coverage. In contrast, a decrease in bare land coverage is observed in the rural area near the mountains. By checking both the urban expansion map and the replacement map, it is worth noting that the mangrove areas in Nadi Bay have gradually been replaced by the urban area since 2016.
Figure 9: Urban expansion map
Figure 10: The replacement map

### v2

@@ Abstract
[inst1]Transitional Artificial Intelligence Research Group, School of Mathematics and Statistics, UNSW Sydney, Sydney, Australia
@@ 2.6.2 Technical setup
In this study, we employed 10-fold cross-validation for model evaluation. The dataset was randomly partitioned into 10 equal subsets, where in each iteration one subset was used for testing and the remaining nine for training. The process was repeated 10 times, and the average performance across all folds was reported. The model achieving the best accuracy among the folds was selected as the final model.
@@ 3.1 Selection of sample size
sample size accuracy precision recall f1 score
490 0.9659 0.9677 0.9659 0.9657
700 0.9523 0.9530 0.9523 0.9521
1050 0.9809 0.9812 0.9809 0.9809
1400 0.9857 0.9881 0.9782 0.9825
1750 0.9943 0.9935 0.9964 0.9949
@@ 3.2 Model construction
Random Forest: The hyperparameters of the random forest algorithm were selected using a five-stage grid search, including ‘max depth’, ‘max features’, ‘min samples leaf’, ‘min samples split’, and ‘n estimators’. The values of each parameter used in the experiment are provided in the table. Model performance was evaluated using 10-fold cross-validation.
Artificial Neural Networks: The model inputs were 1225 flattened vectors derived from image chips of size $9\times 9\times 10$. Hyperparameters were optimized using grid search, including the number of layers, neurons per layer, optimizer, and activation function. Training was monitored on accuracy, with early stopping (patience = 15) to prevent overfitting. The final ANN architecture consisted of one hidden layer with 32 neurons, the ‘Adam’ optimizer, ReLU activation, and categorical cross-entropy as the loss function. Training was conducted for a maximum of 150 epochs with a batch size of 32. Model evaluation was based on 10-fold cross-validation, and the best-performing fold was taken as the final result.
Convolutional Neural Networks: The model inputs were 1225 image chips of size $9\times 9$ pixels with ten channels. The CNN architecture consisted of three convolutional layers with 32, 48, and 64 kernels of size $1\times 1$, followed by dropout layers (0.25 and 0.5) to mitigate overfitting. The output was flattened and passed through a fully connected layer. ReLU activation and the Adam optimizer were used, with sparse categorical cross-entropy as the loss function. Training was performed for up to 150 epochs with early stopping (patience = 15). Model performance was assessed using 10-fold cross-validation, with the best-performing fold taken as the final result.
@@ 3.3 Score comparison
In order to choose the model with the best classification performance, we have compared the accuracy scores of three supervised learning models, which were delivered by dividing the number of correctly classified samples by the total test samples. The classification accuracies for all three models were high. ANN had an accuracy of 98.29%, and CNN had an accuracy of 99.43%. The confusion matrices are shown in the Figure. In general, all three models performed well in classifying the Coastal Areas and wetlands while making mistakes in classifying the Grass/Agricultural Land, Forest and Water Body. We especially cared about the performance of the models in classifying the urban areas. The distribution of houses is sparse in the eastern side of our study area, close to the mountain areas, making distinguishing between bare land and urban area difficult due to their similar spectral reflectance. CNN performed best among the three models. This could be a good sign for us since an accurate classification of the urban areas is essential for change detection.
Table 2: Accuracy of supervised learning models over 10-fold cross-validation.
Fold CNN RF ANN
1 0.9429 0.9714 0.9600
2 0.9657 0.9600 0.9543
3 0.9714 0.9771 0.9657
4 0.9943 0.9771 0.9886
5 0.9829 0.9829 0.9829
6 0.9657 0.9771 0.9714
7 0.9714 0.9543 0.9600
8 0.9771 0.9657 0.9771
9 0.9771 0.9714 0.9771
10 0.9486 0.9486 0.9771
Mean 0.9697 0.9686 0.9714
Std 0.0002090 0.0001123 0.0105374
@@
(a) RF
(b) ANN
(c) CNN
Figure 6: The ROC AUC curves of supervised learning models
@@ 3.5 Landcover Maps
(a) 2014
(b) 2015
(c) 2016
(d) 2017
(e) 2018
(f) 2019
(g) 2020
(h) 2021
(i) 2022
(j) 2023
Figure 8: Classified Images from 2014 to 2023
@@ 3.6 Urban Area change detection
Based on the land cover maps that were generated by CNN, we created an urban expansion map for visualising the change in the urban area. The urban expansion map Figure 9 shows how the urban area was expanded from 2013 to 2023. In the image, the non-urban area is represented by the white color.
The urban expansion map reveals that the urban area in our study has expanded outward along the existing periphery and the growth of the urban area in the Nadi downtown, near the coastal line, corresponds to a substantial reduction in grass and agricultural Land coverage. In contrast, a decrease in bare land coverage is observed in the rural area near the mountains. By checking the urban expansion map, it is worth noting that the mangrove areas in Nadi Bay have gradually been replaced by the urban area since 2016.
Figure 9: Urban Expension Map
@@ 5 Conclusion
The code and dataset are publicly available at our GitHub repository.